a disck initially stationary experiences an angular acceleration of 0.2 rad/s^2. how many revolutions does the disck undergo in the 10 seconds after the acceleration is applied?

a- 0.1

b- 0.80

c- 1.6

d- 8.0

e- 10
Answer #1

Angular acceleration, \alpha=0.2 rad/sec^2

time , t=10 sec

number of revolutions n=(1/2)* \alpha * t^2

                                =0.5*0.2*10*10

                               =10